A merchant has coffee worth $60 a pound that she wishes to miz with 70 pounds of coffee worth $90 a pound to get a mixture that is worth $70 a pound. How many pounds of the $60 coffee should be used?
Found 2 solutions by fcabanski, stanbon:
Answer by fcabanski(1391)

Check my profile and contact me for one on one online tutoring.


We don't know how much $60 coffee should be used. Call it x.


We do know how much of the $90 coffee is used. 70 lbs.


When mixed, there is x + 70 pounds of coffee, and that is worth $70/lb.


We can set up an equation: 60x + 6300 = (x+70)*70.


Solving for x reveals the answer.


60x +6300 = 70x +4900 --> 1400 = 10x ---> 140=x


Answer by stanbon(75887)

A merchant has coffee worth $60 a pound that she wishes to miz with 70 pounds of coffee worth $90 a pound to get a mixture that is worth $70 a pound. How many pounds of the $60 coffee should be used?
----
Equation:
value + value = value
60x + 90*70 = 70*(x+70)
60x + 90*70 = 70x + 70*70
-10x = -20*70
x = 140 (amt. of $60 coffee needed)
===========
Cheers,
Stan H.